The Student Information Processing Board

The SIPB (pronounced sip-bee), or Student Information Processing Board, is a service organization whose purpose is to improve computing at MIT. We’ve been around MIT for over 25 years. Over the years, we’ve introduced services such as the first time-shared computing facilities on campus, and dialup to athena before it was available from the Institute. Today, there are many different services we provide.

SIPB members and prospectives help provide services such as support and maintenance of USENET news, maintenance of MIT’s first world wide web server, and support of various useful software on athena in the sipb locker. We also work on support and development of the Linux and NetBSD versions of Athena and answer user questions on just about any computer-related topic. On the educational side, we run tours of computational facilities at MIT, write help documentation, and teach IAP classes on lots of topics. We also tend to spend a lot of time hanging around the office playing loud music, staying up too late, playing silly computer games, talking, and being friends.

The SIPB office, in w20-557, (next to the Athena Cluster on the fifth floor of the Student Center) is the place where most of SIPB’s activities are based. The office has a number of workstations: PC’s running Linux-Athena and NetBSD-Athena, one of every type of workstation that’s in the public clusters, a few types of workstations which have long since been retired from the clusters, and a Mac which is currently the only publicly available Mac on campus. We’re generally more likely to be open all night than, say, between 9am and noon, although you can often find people still around the office at 9am from the night before. Members and prospectives use the office both for doing productive SIPB work and as a general place to log in and hang out. While hanging out in the office, though, it’s always expected that people are willing to answer the phone, handle walk-in questions, and try not to annoy everyone else in the room.

If you’re interested in the kind of things the SIPB does, whether or not you know a lot about computers already, then you should consider becoming a SIPB prospective. You can become a prospective by coming into the office, finding a SIPB member, and telling them you’re interested in becoming a prospective. Prospectives are people who are interested in becoming a SIPB member, but have not yet been elected as members. Generally, prospectives try to learn about SIPB and learn useful skills for being a SIPB member, by hanging out in the office, talking to people, reading documentation, getting to know the office, and getting to know members. Prospectives also try to attend our weekly meetings at 7:30 pm Monday nights, and generally try to find some useful things to do for SIPB to help out with. After attending at least four meetings, prospectives can be nominated for membership. Becoming a member mostly means that you get a vote on things during meetings, are allowed to know the SIPB root password, and get a key to the office. Usually, it takes longer than a month to become a member, and people often say that the right time to nominate someone for membership is when half of the office forgets that they aren’t already a member.

The next SIPB meeting is next Monday at 7:30pm in the SIPB office, w20-557. We hope to see you at the meeting, or in the office any time, if you would like to become a prospective. You can also reach us by calling x3-7788, or sending mail to sipb@mit.edu. We’re happy to answer your questions about becoming a prospective, or to help you out with any computer-related questions you might have.